# Northeast Materials Database (NEMAD): Enabling Discovery of High Transition Temperature Magnetic Compounds

### v1

@@ Abstract
The discovery of novel magnetic materials with greater operating temperature ranges and optimized performance is essential for advanced applications. Current data-driven approaches are challenging and limited due to the lack of accurate, comprehensive, and feature-rich databases. This study aims to address this challenge by introducing a new approach that uses Large Language Models (LLMs) to create a comprehensive, experiment-based, magnetic materials database named the Northeast Materials Database (NEMAD), which consists of 26,706 magnetic materials (www.nemad.org). The database incorporates chemical composition, magnetic phase transition temperatures, structural details, and magnetic properties. Enabled by NEMAD, machine learning models were developed to classify materials and predict transition temperatures. Our classification model achieved an accuracy of 90% in categorizing materials as ferromagnetic (FM), antiferromagnetic (AFM), and non-magnetic (NM). The regression models predict Curie (Néel) temperature with a coefficient of determination (R2) of 0.86 (0.85) and a mean absolute error (MAE) of 62K (32K). These models identified 62 (19) FM (AFM) candidates with a predicted Curie (Néel) temperature above 500K (100K) from the Materials Project. This work shows the feasibility of combining LLMs for automated data extraction and machine learning models in accelerating the discovery of magnetic materials.
@@ Introduction
The conventional techniques for material discovery have been largely based on systematic exploration of compositional space and intuition-directed experimentation. Although successful, these approaches require a lot of time and resources, and they can take years to produce useful outcomes [7]. The development of computational methods has opened up exciting new avenue toward materials discovery and property prediction. The first principles calculation such as the density functional theory (DFT) has enabled the predictions of material properties[8, 9]. Exploration of vast compositional spaces has been possible using high-throughput computational screening [10, 11]. These computational breakthroughs have resulted in the construction of huge materials databases, such as the Materials Project [10], AFLOW [11], AFLOWLIB [12], and OQMD [13]. Unfortunately, applying DFT to predict the magnetic properties of magnetic materials generally leads to less reliable results. Using the standard exchange-correlation functionals could not accurately describe the strongly correlated electron system in magnetic materials especially itinerant magnets. To accurately describe such a system, one should impose the space and spin symmetry constraints as done in ab initio wave function theory. However, the DFT method could not properly incorporate these constraints[14]. The integration of electronic structure information and mean field models for determining exchange interaction and constructing an effective Hamiltonian is a conventional approach in calculating the Curie temperature[15, 16, 17, 18]. This method, however, receives good results only for a short list of materials[19]. Moreover, the requirement for prior knowledge of the crystal structure of the material also puts a limit on the applicability of the DFT method. Another drawback of DFT calculation is that it is mostly limited to the materials of small unit cells due to the high computational costs of large unit cells[20, 21]. Due to the above constraints, determining the magnetic properties and discovering novel magnetic materials remains a difficult task.
@@
The data-driven approach is a new paradigm of research that has received great interest in recent years because of its capabilities in accurate predictions and discovering hidden patterns that otherwise may get missed by traditional methods[22, 23]. Machine learning predicts physical properties quickly, discovers new materials, and optimizes the existing ones by analyzing large datasets[24, 25]. Specifically, several machine learning models have been developed to classify the magnetic material and predict magnetic properties, including phase transition temperatures and coercivity fields[26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. However, the performance of all these machine learning models highly depends on the availability and quality of a comprehensive magnetic materials database, which has not yet been fully available. It is rather challenging to identify trends across different classes of magnetic materials without such a comprehensive database, potentially overlooking promising research directions. For instance, the models were trained on extremely small data points of high-temperature regions specifically for the phase transition temperature prediction[26, 31]. This may generate less accurate predictions in that region which in turn undermines the discovery of high transition temperature materials. MAGDATA [36, 37] is a manually curated database comprising information for around 2000 magnetic materials with experimentally verified lattice and magnetic structure features. Due to its very modest size of dataset, fewer features, and being designed for single-entry lookups, the implementation of data-driven methods on that is challenging and less effective. Some articles and books have noticed some efforts to create a magnetic material database, but the data are scattered and hard to access electronically and systematically [26, 28, 38, 39, 40].
@@
With the advent of the natural language processing (NLP), its combination with machine learning makes it possible to generate an automated database of magnetic materials[41]. An automatically generated database using the ChemDataExtractor toolkit and Snowball algorithm shows the power of this technique by including 39,822 records. But the database is restricted to only a few features, such as chemical compositions and their associated magnetic phase transition temperatures, which are relatively easy to extract from articles. This approach was less accurate when extracting information involving multiple materials or when the data came in varied sentence structures. Another study has been done to build an auto-generated database of 3,613 ferromagnetic compounds with their corresponding Curie temperature[42]. They used a method based on fine-tuning bidirectional encoder representations from transformers (BERT) models. The database shows high quality and accuracy, making it valuable for developing machine learning models to predict material properties. However, database size and features are limited, and they failed to extract data from the tables of a scientific article. Nevertheless, the machine learning models built on both auto-generated databases showed encouraging performance, as indicated by the statistical parameters R2, MAE, and RMSE.
@@
Figure 1: Workflow for automated magnetic materials database creation and analysis. This work retrieves DOIs of scientific articles, downloads and parses the XML files, and creates a CSV file. Information extraction is handled by the GPTArticleExtractor model, which uses prompt engineering, text tokenization, and FAISS. It produces a consistent JSON output through one-shot prompting. Next, a database is generated and manually validated. This database is then used to develop machine learning models that classify materials as FM, AFM, or NM and predict Curie and Néel temperatures, ultimately screening for high-performance magnetic compounds in the Materials Project database.
In this work, using the GPTArticleExtractor workflow shown in Figure 1, we have developed a database of 26,706 magnetic materials, incorporating chemical composition, structural details (such as crystal structure, lattice structure, lattice parameters, and space group symmetry), and magnetic properties (such as coercivity, magnetization, magnetic moment, remanence, and susceptibility). We then implemented this database to train machine learning models, including Random Forest (RF) classifier, Random Forest (RF) regressor, Extreme Gradient Boosting (XGBoost), and Ensemble Neural Network (ENN) for classifying magnetic materials and predicting transition temperature (Curie and Néel). The performance of these models were evaluated by cross-validation techniques. Our classification model achieved 90% accuracy in classifying materials as ferromagnetic, antiferromagnetic, or non-magnetic. Our best Curie temperature prediction model with features generated from chemical composition (chemical composition + structure ) predicts Curie temperature with an R2 value of 0.86 (0.86) and a mean absolute error of 62K (52K). Similarly, our best Néel temperature prediction model with features generated from only chemical composition predicts Néel temperature with an R2 value of 0.85 and a mean absolute error of 32K. This model identified 62 promising ferromagnetic candidates with predicted Curie temperatures exceeding 500K and 19 antiferromagnetic compounds with predicted Néel temperature greater than 100K.
@@ RESULTS
By applying state-of-the-art LLMs to scholarly experimental articles published in Elsevier journals, see details in the Methods section, we have constructed the NEMAD database to provide complete and accurate information on magnetic materials and their corresponding properties. The database includes 26,706 magnetic materials, uniquely defined by chemical composition, structural detail, and magnetic properties. Each entry contains fifteen features listed in Table 1. The chemical composition of each magnetic material was collected initially as string data, but was later converted into a numerical format by feature engineering techniques detailed in Methods, making it easy for training machine learning models.
@@
Material_Chemical_Composition String Chemical composition of the magnetic compound
@@
Magnetization Numeric A/M Magnetization value of compound
Magnetic Moment Numeric $\mu$B Magnetic moment value of compound
Remanence Numeric A/M Remanence value of compound
@@
NEMAD includes both ferromagnetic and antiferromagnetic materials identified by their temperature record. Around 65% of the records in the database are ferromagnets. Records with solely Néel temperature classified as antiferromagnets make up about 33% of the database. Only a tiny percentage (about 3%) of materials have both Curie and Néel temperatures due to their complicated phase diagrams. The Curie-Weiss temperature is another feature that contains important information for the magnetic properties. Discrepancy between the Curie-Weiss temperature and transition temperature usually indicates competing interactions in the material. The distributions of Curie and Néel temperatures throughout the database are shown in Figures 2(a) and 2(b). It is noteworthy that almost 22% of the compounds have Curie temperatures higher than 600K, suggesting broad temperature distributions and the possibility of using this database to search high Curie temperature materials.
@@
The elemental composition of the NEMAD database is extensive, including 84 different elements. The frequencies of these elements in the database are summarized in Figure 2(c). Materials containing elements with high Curie temperatures, such as iron (Fe), cobalt (Co), and nickel (Ni), are numerous. Furthermore, some other elements like Mn, La, Sr, Cu, B, Al, Cr, and Ce are also present in appreciable frequencies, which are necessary for the construction of high-performance magnetic materials. Noticeably, this database contains a huge number of magnetic compounds without rare earth elements. It is potentially useful for discovering permanent magnets without rare-earth elements. Figure 2(d) displays the distribution of compound types in our database. Approximately 40% are ternary compounds (composed of three distinct elements). The remaining compounds are predominantly quaternary, quinary, and binary.
To validate our work, we randomly chose 200 records from our database and manually reviewed the original article associated with these records. The overall precision of the database is 0.97, with details of the validation process provided in the Supplementary Materials.
@@
To classify the materials into non-magnet (NM), ferromagnet(FM), and antiferromagnet(AFM) categories, we built a random forest classifier model based on NEMAD database, see details in the Methods section. Non-magnetic materials used to train the model are taken from the Materials Project. A consistent accuracy of 0.9 on both validation and testing sets underpins the strong generalization of our model toward unseen data. We provide detailed performance metrics for every class across the training, validation, and testing datasets in Table 2. Precision, recall, and F1- score for all classes on the validation and testing sets are consistent, demonstrating that the model is not biased towards the specific subset of data. The overall performance of the model is slightly better in the training set than the testing set, indicating minor over-fitting but the model still remains robust. The AFM class has a slightly lower value of performance metrics compared to other classes, which may be attributed to the smaller antiferromagnetic records in our training dataset. This conclusion is supported by the consistent under-performance of the AFM class across validation and testing sets. Figure 3(a) and 3(b) present the confusion matrices which further illustrate the classification performance across validation and testing datasets.
Existing classification models in literature usually classify NM, FM, and AFM using two-step approach; the first step is to classify only non-magnetic and magnetic compounds, and only magnetic materials are passed into the second step to classify FM or AFM [32, 30]. Our model achieves one-step classification successfully. Besides, almost all of them are trained on the DFT-generated dataset. In contrast, our model was trained on dataset of first hand experimental reports only. It shows that only input of chemical composition is enough to classify the magnetic materials with appreciable accuracy.
Learning how our models use the input features is key to understand their predictions and the relationships in our data. Machine learning models, especially complex ones like neural networks, are notorious for their black boxes nature; and it is hard to see how they are making predictions based on the input features. Figure 3(c) shows the feature importance score, where the most important features are the average atomic magnetic moment, the average atomic mass, the average atomic electronegativity, and the average atomic number. Other influential features include the presence of Mn, Fe, O, and Co atoms.
Table 2: Evaluation metrics used to assess the performance and effectiveness of classification model.
Dataset Class Precision Recall F1-Score Accuracy
Training (60%) FM 1.00 1.00 1.00 0.99
AFM 0.99 0.98 0.99
@@
Validation (20%) FM 0.92 0.90 0.91 0.90
AFM 0.81 0.75 0.78
NM 0.92 0.97 0.94
Testing (20%) FM 0.90 0.90 0.90 0.90
AFM 0.82 0.76 0.79
NM 0.93 0.96 0.94
(a) Validation Set
(b) Testing Set.
(c) Features Importance plot
Figure 3: Performance and analysis of the classification model for magnetic materials. Confusion matrix illustrates the performance of the classification model on the (a) validation set (b) test set , showing the number of true positive, true negative, false positive, and false negative predictions. (c) The feature importance plot ranks the significance of various features used in the classification model. In the Random Forest model, feature importance scores are computed as the mean and standard deviation of the accumulation of impurity decrease within each tree in the forest. The features are then sorted by these importance scores, and the top 20 most influential features are displayed.
@@
In this section, we developed and compared different regression models, including Random Forest (RF) regressor, Extreme Gradient Boosting (XGBoost), and Ensemble Neural Network (ENN), toward accurate prediction of the Curie temperature for a ferromagnetic compound. These Models were trained on two different datasets, one with all the unique ferromagnetic records in the NEMAD database and another with balanced dataset. This balanced dataset was achieved by randomly removing the data points with lower Curie temperature in the NEMAD database to achieve a more uniform temperature distribution. For details on the dataset distribution, see the model development section in the Methods section. Model performance was measured with standard metrics, including R-squared (R2), mean absolute error (MAE), and root mean squared error (RMSE). Furthermore, 5-fold cross-validation was performed to make the results more generic and robust. Table 3 provides detailed performance metrics of these models for test set of each dataset. This implies that across both datasets, the performance of the XGBoost model was greater than that of RF and the ENN model. For the original dataset, the XGBoost model achieved an R2 value of 0.80, MAE of 66K, and RMSE of 114K, slightly outperforming both RF (R2 = 0.80, MAE = 70K, RMSE = 114K) and ENN (R2 = 0.79, MAE = 63K, RMSE = 118K) . Similarly, for the balanced dataset, the XGBoost model reached an R2 value of 0.86 MAE of 62K, and RMSE of 105K while RF and ENN attained (R2=0.83, MAE = 72K, RMSE = 113K) and (R2=0.84, MAE = 65K, RMSE = 110K) respectively. Notably, all models performed better using the balanced dataset; in particular, performance improved most for the XGBoost model. On this balanced dataset, this resulted in a gain of 6% in predictive accuracy for the XGBoost model and 3% and 5% improvement for the RF and ENN model, respectively. Higher performance on the balanced dataset indicates that models are better at capturing the relationships within higher temperature ranges. This could be an improvement due to the increased relative representation of the higher temperature data.
The plots between the predicted and actual Curie temperature for test set of balanced dataset are shown in Figures 4(d), 4(e), and 4(f). This model based on the balanced dataset has better performance than that on the original dataset, as shown in 4(a), 4(b), and 4(c). In the region where the Curie temperature is greater than 500K in the original dataset, all three models underestimated the Curie temperature. This suggests the necessity of expanding the database by adding higher Curie temperature data points in the future.
To further assess the model’s predictive accuracy, we conducted an error analysis using the absolute error plots. Figure 4(g), 4(h), and 4(i) presents these plots for all three models across balanced dataset. We fitted an exponential curve to each error distribution to characterize the error pattern. In XGBoost and ENN, 63% of the test data have an absolute error of less than 50K, whereas in RF model, this proportion is 55%.
To make the model more understandable we calculated the feature importance by calculating the correlation between the features and the target. Figure 4(j), 4(k), and 4(l) shows the feature importance of top 20 features. In all three models, the proportion of Fe atom, magnetic element proportion, rare element proportion, average atomic magnetic moment, and proportion of Co atom are the top predictors.

### v2

# The Northeast Materials Database for Magnetic Materials

@@ Abstract
The discovery of magnetic materials with high operating temperature ranges and optimized performance is essential for advanced applications. Current data-driven approaches are limited by the lack of accurate, comprehensive, and feature-rich databases. This study aims to address this challenge by using Large Language Models (LLMs) to create a comprehensive, experiment-based, magnetic materials database named the Northeast Materials Database (NEMAD), which consists of 67,573 magnetic materials entries(www.nemad.org). The database incorporates chemical composition, magnetic phase transition temperatures, structural details, and magnetic properties. Enabled by NEMAD, we trained machine learning models to classify materials and predict transition temperatures. Our classification model achieved an accuracy of 90% in categorizing materials as ferromagnetic (FM), antiferromagnetic (AFM), and non-magnetic (NM). The regression models predict Curie (Néel) temperature with a coefficient of determination (R2) of 0.87 (0.83) and a mean absolute error (MAE) of 56K (38K). These models identified 25 (13) FM (AFM) candidates with a predicted Curie (Néel) temperature above 500K (100K) from the Materials Project. This work shows the feasibility of combining LLMs for automated data extraction and machine learning models to accelerate the discovery of magnetic materials.
@@ Introduction
The conventional techniques for material discovery have been largely based on systematic exploration of compositional space and intuition-directed experimentation. Although successful, these approaches require a lot of time and resources, and they can take years to produce useful outcomes [7]. The development of computational methods has opened up exciting new avenue toward materials discovery and property prediction. The first principles calculation such as the density functional theory (DFT) has enabled the predictions of material properties[8, 9]. Exploration of vast compositional spaces has been possible using high-throughput computational screening [10, 11]. These computational breakthroughs have resulted in the construction of huge materials databases, such as the Materials Project [10], AFLOW [11], AFLOWLIB [12], and OQMD [13]. Unfortunately, applying DFT to predict the magnetic properties of magnetic materials generally leads to less reliable results. Using the standard exchange-correlation functionals could not accurately describe the strongly correlated electron system in magnetic materials especially itinerant magnets. To accurately describe such a system, one should impose the space and spin symmetry constraints as done in ab initio wave function theory. However, the DFT method could not properly incorporate these constraints[14]. The integration of electronic structure information with mean-field models is a standard approach for estimating magnetic properties. This methodology is commonly used to determine exchange interactions and construct an effective Hamiltonian when calculating the Curie temperature.[15, 16, 17, 18]. This method, however, receives good results only for a short list of materials[19]. Moreover, the requirement for prior knowledge of the crystal structure of the material also puts a limit on the applicability of the DFT method. Another drawback of DFT calculation is that it is mostly limited to the materials of small unit cells due to the high computational costs of large unit cells[20, 21]. Due to the above constraints, determining the magnetic properties and discovering novel magnetic materials remains a difficult task.
@@
The data-driven approach is a new paradigm of research that has received great interest in recent years because of its capabilities in accurate predictions and discovering hidden patterns that otherwise may get missed by traditional methods[22, 23]. Machine learning predicts physical properties quickly, discovers new materials, and optimizes the existing ones by analyzing large datasets[24, 25, 26]. Specifically, several machine learning models have been developed to classify the magnetic material and predict magnetic properties, including phase transition temperatures and coercivity fields[27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. However, the performance of all these machine learning models highly depends on the availability and quality of a comprehensive magnetic materials database, which has not yet been fully available. It is rather challenging to identify trends across different classes of magnetic materials without such a comprehensive database, potentially overlooking promising research directions. For instance, the models were trained on extremely small data points of high-temperature regions specifically for the phase transition temperature prediction[27, 32]. This may generate less accurate predictions in that region which in turn undermines the discovery of high transition temperature materials. MAGDATA [39, 40] is a manually curated database comprising information for around 2000 magnetic materials with experimentally verified lattice and magnetic structure features. Due to the dataset’s modest size, fewer features and single-entry lookups, the implementation of data-driven methods on that is challenging and less effective. Some articles and books have noticed some efforts to create a magnetic material database, but the data are scattered and hard to access electronically and systematically [27, 29, 41, 42, 43].
@@
With the advent of the natural language processing (NLP), its combination with machine learning makes it possible to generate an automated database of magnetic materials. An automatically generated database using the ChemDataExtractor toolkit and Snowball algorithm shows the power of this technique by including 39,822 records[44]. But the database is restricted to only a few features, such as chemical compositions and their associated magnetic phase transition temperatures, which are relatively easy to extract from articles. This approach was less accurate when extracting information involving multiple materials or when the data came in varied sentence structures. Another study has been done to build an auto-generated database of 3,613 ferromagnetic compounds with their corresponding Curie temperature[45]. They used a method based on fine-tuning bidirectional encoder representations from transformers (BERT) models. The database shows high quality and accuracy, making it valuable for developing machine learning models to predict material properties. However, database size and features are limited, and they failed to extract data from the tables of a scientific article. Nevertheless, the machine learning models built on both auto-generated databases showed encouraging performance, as indicated by the statistical parameters R2, MAE, and RMSE.
@@
In this work, using the workflow shown in Figure 1, we have developed a database of 67,573 magnetic materials, incorporating chemical composition, structural details (such as crystal structure, lattice structure, lattice parameters, and space group symmetry), and magnetic properties (such as coercivity, magnetization, magnetic moment, remanence, and susceptibility). We then implemented this database to train machine learning models, including Random Forest (RF) classifier, XGB Classifier, Random Forest (RF) regressor, Extreme Gradient Boosting (XGBoost), and Ensemble Neural Network (ENN) for classifying magnetic materials and predicting transition temperature (Curie and Néel). The performance of these models were evaluated by cross-validation techniques. Our classification model achieved 90% accuracy in classifying materials as ferromagnetic, antiferromagnetic, or non-magnetic. Our best Curie temperature prediction model with features generated from chemical composition (chemical composition + structure ) predicts Curie temperature with an R2 value of 0.87 (0.83) and a mean absolute error of 56K (52K). Similarly, our best Néel temperature prediction model with features generated from only chemical composition predicts Néel temperature with an R2 value of 0.83 and a mean absolute error of 38K. This model identified 25 promising ferromagnetic candidates with predicted Curie temperatures exceeding 500K and 13 antiferromagnetic compounds with predicted Néel temperature greater than 100K.
Figure 1: Workflow for the Construction and Analysis of a Magnetic Materials Database. Scientific articles are processed via three pathways based on their format. Articles retrieved through the Journal’s API in XML format are parsed using both a text parser and a table parser. Standard PDF documents are handled by a PDF parser, which converts the content into markdown text. For older, scanned or image based PDFs and historical handbooks, we use Google Gemini’s OCR capabilities to extract text and tables accurately. For longer documents like handbooks, the content is processed page-by-page and converted into markdown format. All markdown outputs are then converted into CSV files. These files are passed through GPT-4o with structured prompts to extract relevant materials data in a consistent JSON format. After cleaning and standardizing the extracted information, we compile it into the NEMAD database. The curated dataset is used to train machine learning models for classification and for predicting Curie and Néel temperatures. The trained models are then applied to screen for high-performance magnetic compounds.
@@ Results
By applying state-of-the-art LLMs to scholarly experimental articles published in Elsevier and American Physical Society (APS) journals, see details in the Methods section, we have constructed the NEMAD database to provide complete and accurate information on magnetic materials and their corresponding properties. The database includes 67,573 magnetic materials, uniquely defined by chemical composition, structural detail, and magnetic properties. Each entry contains fifteen features listed in Table 1. The chemical composition of each magnetic material was collected initially as string data, but was later converted into a numerical format by feature engineering techniques detailed in Methods, making it easy for training machine learning models.
@@
Material Chemical Composition String Chemical composition of the magnetic compound
@@
Magnetization Numeric $\mathrm{AM}^{-1}$ Magnetization value of compound
[…]
Remanence Numeric $\mathrm{AM}^{-1}$ Remanence value of compound
@@
NEMAD includes both ferromagnetic and antiferromagnetic materials identified by their temperature record. Around 68% of the records in the database are ferromagnets. Records with solely Néel temperature classified as antiferromagnets make up about 30% of the database. Only a tiny percentage (about 2%) of materials have both Curie and Néel temperatures due to their complicated phase diagrams. The Curie-Weiss temperature is another feature that contains important information for the magnetic properties. Discrepancy between the Curie-Weiss temperature and transition temperature usually indicates competing interactions in the material. The distributions of Curie and Néel temperatures throughout the database are shown in Figure 2. It is noteworthy that almost 22% of the compounds have Curie temperatures higher than 600K, suggesting broad temperature distributions and the possibility of using this database to search high Curie temperature materials.
@@
The elemental composition of the NEMAD database is extensive, including 84 different elements. The frequencies of these elements in the database are summarized in Figure 2c. Materials containing elements with high Curie temperatures, such as iron (Fe), cobalt (Co), and nickel (Ni), are numerous. Furthermore, some other elements like Mn, La, Sr, Cu, B, Al, Cr, and Ce are also present in appreciable frequencies, which are necessary for the construction of high-performance magnetic materials. Noticeably, this database contains a huge number of magnetic compounds without rare earth elements. It is potentially useful for discovering permanent magnets without rare-earth elements. Figure 2d displays the distribution of compound types in our database. Approximately 40% are ternary compounds (composed of three distinct elements). The remaining compounds are predominantly quaternary, quinary, and binary.
To validate the quality of the NEMAD database, we randomly selected 5,015 records and used a another large language model (Google Gemini 2.5) to independently evaluate each extracted field against the corresponding original article. The model assessed each entry based on whether the extracted value matched, was absent, or incorrectly assigned. The overall accuracy of the database is high, with a median accuracy of 94% across all fields. A small number of manually reviewed samples further confirmed the consistency of the validation model. Full validation details and field-level performance metrics are provided in the Supplementary Information.
@@
To classify the materials into non-magnet (NM), ferromagnet (FM), and anti-ferromagnet (AFM) categories, we trained a random forest classifier model based on the NEMAD database; see details in the Methods section. Random Forest was chosen due to its successful application in materials informatics literature, especially on small to medium sized datasets, and its robustness in handling heterogeneous, tabular data. Non-magnetic materials used to train the model are taken from the Materials Project. The classification was performed using features derived from chemical compositions, as described in detail in the Feature Engineering section. A consistent accuracy of 0.89 on validation and 0.90 on testing sets underpins the strong generalization of our model toward unseen data. We provide detailed performance metrics for every class across the training, validation, and testing datasets in Table 2. Precision, recall, and F1-score for all classes on the validation and testing sets are consistent, demonstrating that the model is not biased toward any specific subset of data. The overall performance of the model is slightly better in the training set than in the testing set, indicating minor overfitting, but the model still remains robust. The AFM class has slightly lower performance metrics compared to other classes, which may be attributed to the smaller number of antiferromagnetic records in our training dataset. This conclusion is supported by the consistent underperformance of the AFM class across the validation and testing sets.
Table 2: Evaluation metrics used to assess the performance and effectiveness of classification models.
Model Dataset Class Precision Recall F1-Score Accuracy
Random Forest Classifier Training (60%) FM 1.00 0.99 1.00 1.00
AFM 0.99 0.99 0.99
@@
Validation (20%) FM 0.90 0.91 0.91 0.89
AFM 0.79 0.73 0.76
NM 0.93 0.96 0.95
Testing (20%) FM 0.91 0.91 0.91 0.90
AFM 0.80 0.76 0.78
NM 0.94 0.97 0.95
XGB Classifier Training (60%) FM 1.00 0.99 1.00 1.00
AFM 0.99 0.99 0.99
NM 1.00 1.00 1.00
Validation (20%) FM 0.91 0.92 0.91 0.90
AFM 0.79 0.75 0.77
NM 0.96 0.97 0.97
Testing (20%) FM 0.91 0.92 0.91 0.91
AFM 0.81 0.78 0.80
NM 0.96 0.97 0.96
In addition to the random forest model, we trained an XGB classifier on the same dataset using the same feature set. This model was selected for its track record in structured scientific data and its ability to capture non-linear relationships through gradient boosting. The XGB classifier model demonstrated nearly identical overall performance, with an accuracy of 0.90 on the validation set and 0.91 on the testing set. While there are slight differences in precision, recall, and F1-score for specific classes (FM, AFM, NM), the performance metrics remain highly consistent between the two models. This similarity further confirms the robustness and generalizability of our classification approach. The AFM class again showed slightly reduced precision and recall, consistent with the limited number of AFM training samples.
Figure 3a, 3b, 3d, and 3e present the confusion matrices, which further illustrate the classification performance across the validation and testing datasets. In these matrices, the diagonal elements represent the number of correctly classified samples for each class, while the off-diagonal elements indicate misclassifications. Both classifiers show high accuracy, as most values are concentrated along the diagonal, with relatively few misclassifications between FM, AFM, and NM classes.
Figure 3: Performance and analysis of classification models for magnetic materials. Top row (a-c): Results from the Random Forest (RF) classifier. Bottom row (d-f): Results from the eXtreme Gradient Boosting (XGBoost) classifier. Confusion matrices illustrate the performance on validation sets (a,d) and test sets (b,e), showing the number of true positive, true negative, false positive, and false negative predictions. The feature importance plots (c,f) rank the significance of various features used in each classification model. Feature importance scores are calculated as mean decrease in gini impurity for the RF classifier and as the average gain of splits for the XGB classifier.
Existing classification models in literature usually classify NM, FM, and AFM using two-step approach; the first step is to classify only non-magnetic and magnetic compounds, and only magnetic materials are passed into the second step to classify FM or AFM [33, 31]. Our model achieves one-step classification successfully. Besides, almost all of them are trained on the DFT-generated dataset. In contrast, our model was trained on dataset of first hand experimental reports. It shows that only input of chemical composition is enough to classify the magnetic materials with appreciable accuracy.
Understanding which features contribute most significantly to classification decisions is essential for interpreting model behavior and gaining insight into the underlying feature-property relationships in the dataset. For both the RF classifier and XGB classifiers, we evaluated feature importance to identify the most influential descriptors used during training. Figure 3c and 3f present the top 20 ranked features for the RF and XGB classifier models, respectively. For the RF classifier, feature importance was computed based on the mean decrease in impurity (MDI), which reflects the cumulative reduction in Gini impurity achieved by each feature across all decision trees in the ensemble. The most impactful features include the average atomic weight, average atomic magnetic moment, the proportion of high curie temperature elements, and average electronegativity. For the XGB classifier, we measured feature importance using the average gain. This metric shows how much each feature improves the model’s predictions when it is used to split the data.
@@
In this section, we trained and compared different regression models, including Random Forest (RF) regressor, Extreme Gradient Boosting (XGBoost), and Ensemble Neural Network (ENN), toward accurate prediction of the Curie temperature for a ferromagnetic compound. These models were selected based on their complementary strengths and established success in prior materials informatics research. Tree-based models such as RF and XGBoost are known for their robustness, interpretability, and ability to handle tabular data with complex feature interactions, while neural networks are well-suited for capturing non-linear patterns in the data. These models were trained on two different datasets: one with all the unique experimentally verified ferromagnetic records in the NEMAD database, and another with a balanced dataset. The original dataset exhibits a highly skewed distribution of Curie temperatures, with a significant concentration of materials in the low-temperature regime. This imbalance can lead machine learning models to be biased toward more frequently occurring temperature ranges, degrading their generalization performance for higher-temperature compounds. To address this, we employed a stratified undersampling technique to construct a balanced dataset that provides uniform coverage across the entire Curie temperature range. Specifically, we partitioned the Curie temperature values into discrete bins and applied stratified sampling to ensure equal representation of samples from each bin. This approach enhances the diversity and distributional balance of the training data, helping to mitigate model bias toward overrepresented regions and improving performance across the full temperature spectrum.
The bin-based stratified undersampling was integrated into a cross-validated ensemble training procedure, in which individual models were independently trained on balanced subsets from each fold. A final ensemble consisting of 30 models was trained on the full stratified training set to generate robust predictions and estimate uncertainty. For details on dataset distribution and the full model development pipeline, see the Methods section.
@@
Model performance was measured with standard metrics, including R-squared (R2), mean absolute error (MAE), and root mean squared error (RMSE). Table 3 provides detailed performance metrics of these models for the test set of each dataset. For the original dataset, the XGBoost model achieved an R2 value of 0.84, MAE of 62 K, and RMSE of 107 K, compared to RF (R2 = 0.82, MAE = 71 K, RMSE = 115 K) and ENN (R2 = 0.84, MAE = 60 K, RMSE = 110 K). Similarly, for the balanced dataset, the XGBoost model reached an R2 value of 0.87, MAE of 56 K, and RMSE of 97 K, while RF and ENN attained (R2 = 0.85, MAE = 65 K, RMSE = 104 K) and (R2 = 0.85, MAE = 56 K, RMSE = 103 K), respectively. While the differences are not substantial enough to claim superiority, XGBoost consistently performs at top across multiple metrics and datasets. Notably, all models performed better using the balanced dataset; in particular, performance improved most for the XGBoost model. On this balanced dataset, this resulted in a 4% improvement in R2 for the XGBoost model and 4% and 1% improvement for the RF and ENN models, respectively. Slightly higher performance on the balanced dataset indicates that models are better at capturing the relationships within higher temperature ranges. This could be an improvement due to the increased relative representation of the higher temperature data.